# A Survey of Language-Based Communication in Robotics

### v1

@@ 5.3.2. Trust, Responsibility, and Governance
A complete study of trust in LLMs is beyond the scope of this work, for a more in-depth discussion we refer the reader to the work of Sun et al. (Sun et al., 2024). The architecture of communication in LLM-based systems we describe goes beyond the simple “dyadic” (1 human and 1 agent) style of interaction that is often studied and instead introduces a complex relationship of trust between different parts of the total collective (Schneiders et al., 2022). In these circumstances the specification of trust becomes more challenging and relies on a variety of fields and ideas (Abeywickrama et al., 2023). A core issue in all aspects of autonomous system deployment, especially with LLMs is that their apparent ability can engender a very high level of trust from a user; while this may seem positive, it can cause people to rely on the system without checking its outputs. Mitigations for this problem do exist, but are so far limited to rephrasing the agent’s outputs slightly (Kim et al., 2024b). There is also a concern around “deskilling” of operators of these systems who rely on effective systems and gradually lose their own abilities to verify the accuracy of the outputs (Choudhury and Chaudhry, 2024). Some have proposed that the outputs of an LLM should be empirically verified instead of checked by a human, which would avoid the over-trust issue, although this is not always possible especially in robotics (Zhou et al., 2024).
@@
Ideally, a model should be able to match its output with an appropriate level of confidence. A model which can do this reliably is said to be well calibrated; calibration in LLMs is hard to quantify but is generally quite poor (Spiess et al., 2024). This is especially true for LLM-generated code which is a popular option in the robotics domain, making trust calibration an important problem (Virk et al., 2024). The increasing prevalence of LLM-generated data online also raises the concern of a “self-referential learning loop” where agents learn from other AI-generated content (Choudhury and Chaudhry, 2024). The implications of this are yet to be seen but may include a loss of context or availability bias in the training data of future systems (Knapton, 2023).
@@
There are also many ethical and legal issues surrounding the liberal and irresponsible use of LLMs, especially in domains such as healthcare where human lives are by definition always at stake (Ong et al., 2024; Meskó and Topol, 2023). This is even now becoming an issue in the legal domain, with an apparent need for practitioners to make themselves aware of the limitations and pitfalls of these models (Roberts Jr., 2023). There is an additional concern of who is responsible for the inevitable mistakes and potential harms caused by this technology; in a robotics context this could cause severe loss of life and it not always apparent whether the fault lies with the designer of the robot, the LLM, or the deployer of the whole system into a working environment is responsible (Yazdanpanah et al., 2023). With LLMs becoming renowned for their abilities there is an apparent asymmetry between credit for positive outcomes and blame for negative ones (Porsdam Mann et al., 2023)
@@
As an emerging technology LLMs face substantial obstacles in terms of gaining an appropriate and balanced level of trust from the public, forming part of a sensible allocation of responsibility, and being well regulated and understood by the various stakeholders involved. When integrated into robots, this potentially becomes even more important, as robots can be dangerous when operating alongside humans. For language-based robots to be feasible there is need for significant progress in the understanding of these issues.
6. Conclusion
@@
As Language Models continue to improve and are integrated into a wider variety of computer systems, it is necessary to consider their utility and impact in the domain of robotics. Language can act as an input space for models which, through breakdown or generation of code and rewards, conveniently allow untrained human operators to command robots in the same way they would with human colleagues. This has ramifications for collaborate robotics and may facilitate increasing prevalence and practicality of autonomous systems in people’s home and work environments. By the same logic, robots can use language as a means to communicate their actions and intentions to humans, as well as actively using language as a bidirectional interface for clarifications. This allows the twin advantages of humans and robots to be combined in an efficient way, again without need for the human to learn a programming language or other technical details.
@@
Extending this, language can be used as the communication space for groups of agents to coordinate and organise. Much like in human groups, language is generally a common means of communication that allows differently trained and aligned platforms to disseminate information as well as building complex collective plans. Finally, language and language-orientated models can be used as the fundamental basis for the architecture of powerful models. This permits an overall information pipeline between multiple humans and multiple robots where any or every component part could feasibly operate entirely on language, allowing human participation and oversight across the entire observation, decision, communication, and action process.
@@
Language models are inherently limited in several ways. They suffer from hallucination and lack a true understanding or model of the environment in which they operate. Additionally, both safety and security make robotics applications ethically and financially risky. These areas are actively being investigated and some approaches are promising. It is likely that language can never practically be used for the entire control flow of a robotic system. Rather, select parts of the whole can be language-based, with more reliable and tested approaches being used where required. Language is a fundamentally irreplaceable ingredient of human intelligence, both social and individual, making it an exciting and necessary avenue for investigation in robotics and likely a cornerstone of many future systems.
Acknowledgements.

### v3

@@ 5.3.2 Trust, Responsibility, and Governance
A complete study of trust in LLMs is beyond the scope of this work, for a more in-depth discussion we refer the reader to the work of Sun et al., (2024). The architecture of communication in LLM-based systems we describe goes beyond the simple “dyadic” (1 human and 1 agent) style of interaction that is often studied and instead introduces a complex relationship of trust between different parts of the total collective Schneiders et al., (2022). In these circumstances, the specification of trust becomes more challenging and relies on a variety of fields and ideas Abeywickrama et al., (2023). A core issue in all aspects of autonomous system deployment, especially with LLMs is that their apparent ability can engender a very high level of trust from a user; while this may seem positive, it can cause people to rely on the system without checking its outputs. Mitigations for this problem do exist but are so far limited to rephrasing the agent’s outputs slightly Kim et al., 2024a . There is also a concern around the “deskilling” of operators of these systems who rely on effective systems and gradually lose their own abilities to verify the accuracy of the outputs Choudhury and Chaudhry, (2024). Some have proposed that the outputs of an LLM should be empirically verified instead of checked by a human, which would avoid the over-trust issue, although this is not always possible, especially in robotics Zhou et al., (2024).
@@
Ideally, a model should be able to match its output with an appropriate level of confidence. A model which can do this reliably is said to be well calibrated; calibration in LLMs is hard to quantify but is generally quite poor Spiess et al., (2024). This is especially true for LLM-generated code which is a popular option in the robotics domain, making trust calibration an important problem Virk et al., (2024). The increasing prevalence of LLM-generated data online also raises the concern of a “self-referential learning loop” where agents learn from other AI-generated content Choudhury and Chaudhry, (2024). The implications of this are yet to be seen but may include a loss of context or availability bias in the training data of future systems Knapton, (2023).
@@
There are also many ethical and legal issues surrounding the liberal and irresponsible use of LLMs, especially in domains such as healthcare where human lives are by definition always at stake Ong et al., (2024); Meskó and Topol, (2023). This is even now becoming an issue in the legal domain, with an apparent need for practitioners to make themselves aware of the limitations and pitfalls of these models Roberts Jr., (2023). There is an additional concern about who is responsible for the inevitable mistakes and potential harms caused by this technology; in a robotics context this could cause severe loss of life and it is not always apparent whether the responsibility lies with the designer of the robot, the LLM, or the deployer of the whole system into a working environment Yazdanpanah et al., (2023). With LLMs becoming renowned for their abilities there is an apparent asymmetry between credit for positive outcomes and blame for negative ones Porsdam Mann et al., (2023).
@@
As an emerging technology, LLMs face substantial obstacles in terms of gaining an appropriate and balanced level of trust from the public, forming part of a sensible allocation of responsibility, and being well-regulated and understood by the various stakeholders involved. When integrated into robots, this potentially becomes even more important, as robots can be dangerous when operating alongside humans. For language-based robots to be feasible there is a need for significant progress in the understanding of these issues.
6 Conclusion
@@
As Language Models continue to improve and are integrated into a wider variety of computer systems, it is necessary to consider their utility and impact in the domain of robotics. Language can act as an input space for models which, through breakdown or generation of code and rewards, conveniently allow untrained human operators to command robots in the same way they would with human colleagues. This has ramifications for collaborative robotics and may facilitate the increasing prevalence and practicality of autonomous systems in people’s home and work environments. By the same logic, robots can use language as a means to communicate their actions and intentions to humans, as well as actively using language as a bidirectional interface for clarifications. This allows the twin advantages of humans and robots to be combined in an efficient way, again without the need for the human to learn a programming language or other technical details.
@@
Extending this, language can be used as the communication space for groups of agents to coordinate and organise. Much like in human groups, language is generally a common means of communication that allows differently trained and aligned platforms to disseminate information as well as build complex collective plans. Finally, language and language-orientated models can be used as the fundamental basis for the architecture of powerful models. This permits an overall information pipeline between multiple humans and multiple robots where any or every component part could feasibly operate entirely on language, allowing human participation and oversight across the entire observation, decision, communication, and action process.
@@
Language models are inherently limited in several ways. They suffer from hallucinations and lack a true understanding or model of the environment in which they operate. Additionally, both safety and security make robotics applications ethically and financially risky. These areas are actively being investigated and some approaches are promising. It is likely that language can never practically be used for the entire control flow of a robotic system. Rather, select parts of the whole can be language-based, with more reliable and tested approaches being used where required. Language is a fundamentally irreplaceable ingredient of human intelligence, both social and individual, making it an exciting and necessary avenue for investigation in robotics and likely a cornerstone of many future systems.
Acknowledgments